OpenStudy (anonymous):

Of the five letters (A,B,C,D,E), two unique letters are to be selected at random. How many possible selections? Please explain.

iYuko (iyuko):

Something like that

jimthompson5910 (jim_thompson5910):

For slot 1, you have 5 choices After you pick a letter, you have 4 choices for slot 2. If order mattered, then you would just multiply 5*4 = 20 But order doesn't matter in this case (AB is the same as BA), so you divide by 2 to correct for double counting 20/2 = 10 so that's why the answer is 10

iYuko (iyuko):

= 5x4/2x1 = 10 there are 10 possible selections

jimthompson5910 (jim_thompson5910):

iYuko that first method is flawed unfortunately. If you tried it with say six letters, you would get 6*2 = 12 but if you had 6 letters, the actual answer is (6*5)/2 = 30/2 = 15

OpenStudy (anonymous):

Thank you! it makes sense now

jimthompson5910 (jim_thompson5910):

yw